a motorboat traveled for 2 hours with an 8 km/h current. the return trip against the same current took 3 hours. find the speed of the motorboat in still water.
Answer by xcentaur(357)

let the speed of the motorboat be x kmph
the speed of the current be y kmph = 8 kmph


now,journey with the current:
speed=x+y=(x+8)kmph,time=2 hours
d=2(x+8)=(2x+16)km


now considering return journey:
speed=x-y=(x-8)kmph,time=3 hours
distance=3(x-8)=(3x-24)km


but the distance in both cases must be the same,
hence 2x+16=3x-24
16+24=3x-2x
x=40 kmph


Thus speed of the motorboat in still water=40kmph


Hope this helps,
good luck.